OpenStudy (anonymous):

600 J of heat is added to a monatomic gas in a process in which the gas performs a work of 150 J . The molar heat capacity for the process is ? The answer is 2R . Please HELP !

OpenStudy (anonymous):

\[Q=U+W . \] Therefore , U= 600-150 =450 J . Molar heat capacity is defined as the amount of energy needed to raise the temperature of 1 mole of substance by 1 Kelvin . \[C = dQ/(dT)n\]

OpenStudy (anonymous):